OpenStudy (anonymous):

Explain why 1 raised to any power is equal to 1 ?

OpenStudy (anonymous):

1^n , means one multiplied n times , however small or large the number be , 1^ any power is always equal to one , even if you consider negative number

OpenStudy (anonymous):

what about \(1^{\sqrt2}\) ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

yes i know, but it does not conform to your reasoning

OpenStudy (anonymous):

\[n \in R\]

OpenStudy (anonymous):

since exponentiation is not repeated multiplication as multiplication is not repeated addition just being obnoxious, ignore me

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i agree, but it is not because "1^n , means one multiplied n times "

OpenStudy (anonymous):

ooh , then why is it

OpenStudy (anonymous):

that reasoning doesn't even work for \(1^0\)

OpenStudy (anonymous):

as you cannot multiply a number by itself zero times

OpenStudy (anonymous):

\[\huge x \in R -0\]

OpenStudy (anonymous):

Is it the correct domain

OpenStudy (anonymous):

i have no idea what that means what about \(1^{-1}\) or \(1^{\pi}\)

OpenStudy (anonymous):



OpenStudy (anonymous):

"1^n , means one multiplied n times " \(\large 1^{\pi}\) means one multiplied \(\pi\) times "?? no meaning there

OpenStudy (anonymous):

"1^n , means one multiplied n times " \(1^{-1}\) means one multiplied \(-1\) times " ?? nope

OpenStudy (anonymous):

2^n , means 2 multiplied n times

OpenStudy (anonymous):

2^3 = 2*2*2

OpenStudy (anonymous):

1^3 = 1*1*1

OpenStudy (anonymous):

\(2^{\sqrt2}\) means \(2\) multiplied by itself \(\sqrt2\) times lol

OpenStudy (anonymous):

your reasoning only work for \(n\in \mathbb{N}\) not for \(n\in \mathbb{Z}\) or \(n\in\mathbb{Q}\) or \(n\in \mathbb {R}\)

OpenStudy (anonymous):

FOR 1^-1 You obviously know this we use indices property It is weird to imagine that but it is

OpenStudy (anonymous):

Yeah , i know , but the dude asking the question , i am presuming he won't all this

OpenStudy (anonymous):

@Hahuja Do you also wanta a explaination for why n^0 = 1?

OpenStudy (anonymous):

@Adjax 1) by definition ( a not very satisfactory but true answer) unless 2) you can precisely define \(b^x\) for all \(x\in \mathbb{R}\) \(b>0\)

OpenStudy (anonymous):

\[\huge 2^{\sqrt{2}}\] means 2 multiplied by root 2 times

OpenStudy (anonymous):

How to multiply a number by root 2 times?

OpenStudy (anonymous):

you cannot

OpenStudy (anonymous):

yeah you cannot

OpenStudy (anonymous):

but its what it is

OpenStudy (anonymous):

@satellite73 ..thats all pure math stuff(a college lvl. stuff)..and here the frame of reference is not that high enough

OpenStudy (anonymous):

for that matter you cannot even multiply something by itself zero times or two thirds times or minus five times

OpenStudy (anonymous):

there's a very good explaination on betterexplained.com by mr.Kalid Azad on this matter from the point of beginner 's view

OpenStudy (anonymous):

that is probably true, you resort to "because it is" or "by definition" or "it seems obvious that"...

OpenStudy (anonymous):

If you can't how do you calculate it what's the proof

OpenStudy (anonymous):

if you want to see why \(2^0\) or \(b^0\) ought to be \(1\) then look at the pattern \[2^4=16\\ 2^3=8\\ 2^2=4\\ 2^1=2\\ 2^0=?\] pretty clear you are diving by \(2\) each time

OpenStudy (anonymous):

'what's the proof": Pursue a degree in pure mathematics/...that will give answer to all yar questions and that will give you a handful of greek letters that why a slice of pizza exists(book way:why a fraction/fractional no. exist?)

OpenStudy (anonymous):

that does not prove that it is 1 , but it suggests that 2^0 might be 1

OpenStudy (anonymous):

What is the course to take to solve such problems?

OpenStudy (anonymous):

here's the link: http://betterexplained.com/articles/understanding-exponents-why-does-00-1/

OpenStudy (anonymous):

if you want to know why \(2^0=1\) other than by definition, what you need to know is that \[b^x:=e^{x\log(b)}\]

OpenStudy (anonymous):

@satellite73 : your above e.h. is not a kind of..err...what about a different no. n instead of 0.. one more link: http://betterexplained.com/articles/an-intuitive-guide-to-exponential-functions-e/

OpenStudy (anonymous):

Is the log there the natural log?

OpenStudy (anonymous):

there is only one log, but yes

OpenStudy (anonymous):

well , to real numbers power its true try to learn complex :D u will start to see things

OpenStudy (anonymous):

and yes your above eqn. reminded me of a EQN> in my fogged memory which proved that for a complex no. Z to the power of real no. N can be evaluated

OpenStudy (anonymous):

a number different from zero is just a different number \[2^\sqrt2=e^{\sqrt2\log(2)}\]

OpenStudy (anonymous):

why man invented zero?

OpenStudy (anonymous):

REal nos. are subset of complex no..